Circular No. 8194

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF URANUS M. R. Showalter, Stanford University; and J. J. Lissauer, Ames Re-

search Center, NASA, report the recovery of two inner Uranian satellites, S/1986 U 10 (cf. IAUC 7171) and Uranus VII (Ophelia), in recent data from the High Resolution Channel of the Hubble Space Telescope (HST) Advanced Camera for Surveys. Images were obtained during three consecutive HST orbits between Aug. 25.8042 and 25.9659 UT. This is the first clear groundbased detection of either satellite and substantially reduces uncertainties in each body’s orbital mean motion. Each satellite is visible as a streak in two consecutive coadded images, each consisting of four 250-s clear-filter (300–900 nm) exposures. Uranus VII crossed the southern ansa of Uranus’s rings at Aug. 25.8819, trailing its predicted orbital longitude by 62^{o}.8 ± 1^{o}.3. This slightly exceeds the 58^{o} uncertainty in its extrapolated orbit; the implied change in its mean motion is −0^{o}.00977 day^{−1}, to 956^{o}.41856 ± 0^{o}.00020 day^{−1}. Uranus VII is ≈ 15 percent as bright as Uranus XIV (Belinda) in the images, with signal-to-noise ratio (S/N) > 20. S/1986 U 10 crossed the northern ansa of the rings at Aug. 25.8264, leading its predicted orbital longitude by 47^{o}.6 ± 0^{o}.9. This is consistent with the 64^{o} uncertainty in its extrapolated orbit; the implied change in its mean motion is +0^{o}.00741, to 564^{o}.24741 ± 0^{o}.00014 day^{−1}. S/1986 U 10 is ≈ 8 percent as bright as Uranus XIV, with S/N > 10.

COMET C/2002 VQ_{94} (LINEAR) An apparently asteroidal object reported by LINEAR (announced on

MPEC 2002-V71, where B. G. Marsden noted “whether this object is a comet or not is inconclusive”, and MPS 66506; discovery observation given below) has been found to have a prominent 10′′ coma with a fanlike morphology spanning p.a. 180^{o}–300^{o} on images taken by D. Jewitt on Aug. 28.5 UT with the University of Hawaii 2.2-m telescope. Recent astrometry, the orbital elements below, and an ephemeris appear on MPEC 2003-R22.

2002 UT  α_{2000}  δ_{2000}  Mag.
Nov. 11.23766  2^{h}02^{m}44.^{s}75  +46^{o}00′35′′.6  18.7

Epoch = 2006 Jan. 25.0 TT

T = 2006 Feb. 6.6544 TT  ω = 100.^{o}0365
e = 0.966862  Ω = 35.0164 }2000.0
q = 6.796699 AU  i = 70.5155
2003 September 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green